# LACONIC: A 3D Layout Adapter for Controllable Image Creation

### v1

@@ A.4.2 Diffusion Model from Scratch
An important design choice of our framework is to rely on a cross-attention-based adapter network, in order to augment existing T2I models with 3D layout guidance, instead of training dedicated generative models from scratch. The main intuitions behind this critical choice are that (i) considering the small scale of available image datasets featuring 3D layout annotations, achieving photorealistic generation results would be extremely challenging, especially on challenging camera views that are less represented in the training distribution and (ii) we could leverage prior knowledge from other domains to achieve powerful generalization, as shown throughout our experimental results in the main paper, which would not be possible otherwise. We believe that those insights validate the relevance of this baseline approach in the context of our contributions.
@@
To experimentally validate our intuitions, we propose to train jointly a randomly initialized diffusion UNet backbone and the 3D layout encoder detailed in Section A.1.2. The backbone is similar to the one of Stable Diffusion, implemented with Diffusers [66], and is guided through regular cross-attention conditioning [53] between the 3D scene sequence $\mathcal{S}$ and feature maps from the residual blocks. Considering the very little variety of image captions in the training datasets, it is not trained with additional text conditioning. The UNet is also downsized, compared to Stable Diffusion, to account for the smaller scale of the training set. Importantly, we stack $4$ downsampling and upsampling residual blocks with $128$, $256$, $256$ and $384$ output channels. The $2$ bottleneck blocks, that operate at smaller resolutions, are augmented with cross-attention with $8$ heads. For the 3D layout encoder, we keep the same parameters as those described in Section A.1.2.
A.5 Evaluation Metric
In Section 4.1 of our main paper, we introduce the Scene Object CLIP (SOC) score as a robust way to simultaneously assess that objects in the generated content (i) are correctly positioned and sized with respect to their spatial conditioning information and, at the same time, (ii) match their assigned semantic attributes.
@@ A.5 Evaluation Metric
To do so, given a conditioning semantic 3D layout $\mathcal{S}$ and camera pose $\mathcal{C}$, its associated ground truth image $x_{0}$ with 2D object $o_{i}$ bounding box annotations $\mathcal{B}_{\text{2D}}^{i}$ and, importantly, a synthesized image that we wish to evaluate, we identify the main visible objects composing the rendered scene. As 2D bounding boxes $\mathcal{B}_{\text{2D}}^{i}$ from the ground truth annotations alone do not assess that the corresponding objects $o_{i}$ are visible in their most part, we also leverage the 3D bounding box from the conditioning signal projected to image coordinate, from which we derive the enclosing 2D bounding box $\mathcal{B}_{\text{3D}}^{i}$. We notice that this second bounding fully contains the one from the image annotation, and compute the ratio $r_{a}$ of the area of $\mathcal{B}_{\text{2D}}^{i}$ over the one of the enclosing projected box $\mathcal{B}_{\text{3D}}^{i}$. Remarkably, this projection mechanism also handles objects that partially lie outside the image bounds. Based on this ratio, objects $o_{i}$ that are sufficiently visible are cropped at the location of $\mathcal{B}_{\text{3D}}^{i}$ in the synthesized image, yielding per-object images $x_{o_{i}}$. We finally compute a CLIP [51] correlation score between object images and their associated text semantic $s_{i}$.
@@
We rely on an object’s associated visible area ratio $r_{a}^{i}$ to determine if its crop image will be added to the evaluation base. In practice and to report the values in Table 1 in the main paper, we set a threshold value $\alpha=0.4$. Also note that tiny objects, that are widely represented in the HyperSim [52] dataset are initially filtered out. Indeed, objects whose instance semantic map spans less than $2$% of the image pixels are discarded. Finally, we also filter objects whose NYU40 semantic annotation is not meaningful, e.g., that are classified as ”other”. The metric is averaged over all the resulting object crops in the dataset.
Appendix B Additional Results
@@
We provide additional qualitative results on text-driven, layout-guided image synthesis in Figure LABEL:fig:teaser. This highlights the ability of our method to adhere with the input semantic 3D layout while being capable of out-of-distribution generalizations to various unseen concepts. Additionally, we report in Figure 9 and Figure 10 supplemental comparisons between the DM-FS and SceneCraft [73] baselines, both with and without providing an input text as an additional condition. Our method consistently generates more realistic images, with better details as well as 3D layout and text prompt adherence. Motivating our 3D layout conditioning encoder, the DM-FS baseline appears to better respect the input 3D layout in comparison to SceneCraft [73]. We notice that both adapter-based approaches, ours and SceneCraft, leverage the pretrained T2I backbone, resulting in less artifacts in the generated images in comparison to DM-FS. Consequently, our method also show strong competitive advantages over SceneCraft and DM-FS when the number of objects in the scene is high thanks to the combination of the expressiveness of our 3D layout conditioning encoder coupled with the ability of the T2I backbone to generate detailed visual features.
Appendix C Limitations
@@
Figure 8: Failure Cases and Limitations. In (a), some objects, such as the windows, are missing in the generated image and the floor shape, while complex, is not accurately respected. We can also observe visual stability issues, as shown in the result in (b), which features an inconsistent pattern on the wall resembling a misplaced frame above the bed. Finally, generated images can exhibit unnatural perspective, resulting in distorted floors and objects, as showcased in (c). This behavior may result from our assumption of consistent camera intrinsics across all data samples.
We provide failure cases and known issues in Figure 8. While showcasing state-of-the-art results on 3D layout adherence, our method can still lead to some inconsistencies in the generated results.
Appendix D Societal Impact
@@
Controllable generative AI for 3D environments has significant implications for both individuals and industries. We believe our method to predominantly yield positive societal impact. By enabling fine-grained control over generation, our approach makes 3D-aware content creation more intuitive and accessible, empowering not only artists and designers but also individuals with no prior expertise in 3D modeling. It also benefits industries that rely on realistic environmental rendering, such as urban planning, architecture, and digital twins for cities. A key advantage of our method is its parameter efficiency, which reduces the computational cost of training—an important factor given the high environmental impact of generative models [33, 27]. However, these benefits come with challenges. Since our approach does not natively incorporate recent advances in generative model watermarking [10, 57], it could be misused for potentially deceptive applications. Additionally, by significantly lowering the barrier to content generation, it may contribute to rapid job displacement in traditional 3D design fields. Another concern is the potential for bias propagation, as image generative models, including semantic and style-driven ones like ours, may unintentionally reinforce stereotypes [80]—for example, by associating certain colors with gender, by e.g., synthesizing a bedroom as blue for ”boys” and pink for ”girls”. Ensuring the responsible development and use of such technologies, while addressing ethical concerns and mitigating biases, is crucial.
@@
Figure 9: Additional 3D layout-guided image synthesis baseline comparisons (1/2). We can observe that our method produces more natural images and with respect to the input 3D layout.

### v2

@@ A.5.2 Diffusion Model from Scratch
A key design choice in our framework is to use a cross-attention-based adapter network to augment existing T2I models with 3D layout guidance, rather than training a dedicated, conditional generative model from scratch. This choice is motivated by two primary factors. First, achieving photorealistic generation results would be extremely challenging given the small scale of available image datasets featuring 3D layout annotations, especially on less common camera views that are not widely represented in the training distribution. Second, our adapter approach allows us to leverage the powerful priors of large-scale T2I models, enabling strong generalization to other domains, as demonstrated by our experimental results. We believe that those insights validate the relevance of this baseline in the context of our contributions’ evaluation.
@@
To experimentally validate our intuitions, we trained a baseline from scratch, jointly optimizing a randomly initialized UNet and the 3D layout encoder from Section A.1.2. The backbone is similar to that of Stable Diffusion, implemented with Diffusers [70], and is guided through standard cross-attention conditioning [55] between the 3D scene sequence $\mathcal{S}$ and feature maps from the residual blocks. Due to the limited caption variety in the training data, we omit text conditioning for this baseline. To account for the smaller training set, the UNet is also downsized compared to Stable Diffusion and comprises 444 downsampling and upsampling residual blocks with 128128128, 256256256, 256256256 and 384384384 output channels. The two bottleneck blocks, that operate at smaller resolutions, are augmented with cross-attention with 888 heads. The 3D layout encoder’s architecture is identical to that described in Section A.1.2.
Figure 9: Layout-guided image synthesis results with a DiT-based backbone. Our LACONIC adapter successfully conditions Stable Diffusion 3 [8], demonstrating compatibility with modern DiT architectures. The generated images (Right) show strong adherence to the input 3D layouts and camera viewpoints (Left).
A.6 Evaluation Metric
In Section 4.1 of our main paper, we introduce the Scene Object CLIP (SOC) score as a robust way to simultaneously assess whether objects in the generated content (i) are correctly positioned and sized with respect to their spatial conditioning information and, at the same time, (ii) match their assigned semantic attributes.
@@ A.6 Evaluation Metric
To compute the score for a synthesized image, we first identify its main visible objects given the conditioning 3D layout $\mathcal{S}$, the camera pose $\mathcal{C}$, and the 2D object $o_{i}$ bounding box annotations $\mathcal{B}_{\text{2D}}^{i}$ from the associated ground truth image $x_{0}$. Since these 2D boxes do not assess that the corresponding objects $o_{i}$ are mostly visible, we also project the 3D bounding boxes from the conditioning signal onto the image plane to derive enclosing 2D boxes $\mathcal{B}_{\text{3D}}^{i}$. For each object, we notice that this second bounding box fully contains the one from the image annotation, and compute the ratio $r_{a}$ as the area of $\mathcal{B}_{\text{2D}}^{i}$ divided by the area of the enclosing projected box $\mathcal{B}_{\text{3D}}^{i}$. Remarkably, this projection mechanism also handles objects that partially lie outside the image bounds. Based on this ratio, objects $o_{i}$ that are sufficiently visible are cropped at the location of $\mathcal{B}_{\text{3D}}^{i}$ in the synthesized image, yielding per-object images $x_{o_{i}}$. We finally compute a CLIP [53] correlation score between object images and their associated text semantic $s_{i}$.
@@
We rely on an object’s associated visible area ratio $r_{a}^{i}$ to determine if its crop image will be added to the evaluation base. In practice and to report the values in Table 1 in the main paper, we set a threshold value $\alpha=0.4$. Additionally, we filter out tiny objects, which are prevalent in the HyperSim [54] dataset. Indeed, objects whose instance semantic map covers less than 222% of the image area are discarded. Finally, we also filter objects whose NYU40 semantic annotation is non-descriptive, e.g., that are classified as “other”. The final metric is averaged over all the resulting object crops in the dataset.
Appendix B Generalization to DiT Backbones
In this section, we demonstrate the flexibility of our adapter by integrating it with a recent Diffusion Transformer (DiT) [48] backbone. Specifically, we employ Stable Diffusion 3 [8] to showcase that LACONIC is not limited to established UNet-based models conditioned via standard cross-attention, but is capable of remarkable generalization across architectures.
B.1 Background on Joint-Attention
Stable Diffusion 3 [8] introduces a Multimodal Diffusion Transformer (MM-DiT) architecture trained with Rectified Flows [29]. The core of its conditioning approach is a joint-attention mechanism that operates over a unified sequence containing both image tokens (from the latent $x_{t}$) and text tokens (from the caption $c$). Within each transformer block, separate sets of linear projections are used to yield the respective image $Q^{x}$, $K^{x}$, $V^{x}$ and text $Q^{c}$, $K^{c}$, $V^{c}$ matrices. These are subsequently concatenated into global matrices (e.g., $Q=[Q^{x};Q^{c}]$) used for attention, enabling a bidirectional information flow between text and visual features within the architecture. The resulting hidden states $H^{x}$ and $H^{c}$ are finally separated and processed by additional modality-specific feedforward modules.
Figure 10: Qualitative comparison with Build-A-Scene [7] given a common 3D layout and viewpoint. For each method, from left to right: input scene representation, generation result for prompt $c_{1}=$a cozy bedroom with a wooden floor”, and for $c_{2}=$a Van-Gogh style bedroom”. Build-A-Scene’s input representation also features individual prompt for each bounding box.
B.2 Layout Conditioning
Our 3D layout condition is integrated by performing an additional joint-attention pass. Consistent with our general methodology, this is implemented by introducing new trainable linear projections to augment the pretrained model. Specifically, we obtain new key-value pairs from both the text-conditioned image state $H^{x}$ and the encoded layout $y$ features, while leveraging the query matrix $Q$ from the primary text-conditioning step. A second joint-attention operation is then performed over concatenated matrices, producing an output that is split into a novel image state $H^{x^{\prime}}$, and a layout state $H^{y}$ which is processed by additional linear layers. These are merged with the outputs from the primary joint-attention pass, following the linear combination from Equation (3) in the main paper, i.e.,
$H^{x}_{\text{final}}=H^{x}+\gamma H^{x^{\prime}}\quad\text{and}\quad H^{c}_{\text{final}}=H^{c}+\gamma H^{y}$ (6)
In practice, we leverage the Diffusers backbone implementation and pretrained weights, that we augment with the custom joint-attention processor implementation from IPAdapter-Instruct [57].
B.3 Qualitative Results
We show in Figure 9 images synthesized by Stable Diffusion 3 [8] DiT after training a dedicated LACONIC adapter. Notably, the augmented backbone successfully takes into account the input 3D layout and viewpoint and, as a result, renders scenes in accordance with the specified objects spatial and semantic features. Additionally, Figure 8 highlights that the model also supports multimodal conditioning with a non-empty global text caption $c$, whose influence with respect to the layout can be adjusted following Equation (6).
Appendix C Additional Results
In this section, we provide additional experimental results highlighting the capabilities of our model and advantages of our design choices.
C.1 Ablation Study
We trained ablated versions of our model to measure the individual contribution of:
  • L-R: applying the reframing mechanism from world to camera coordinates described in Section 3.4.
  • L-T: including the transformer encoder module in the adapter architecture detailed in Section 3.3.
Quantitative evaluation metrics are computed on the custom bedroom dataset and reported in Table 2. Reported results further validate our design choices. Notably, we can observe that the transformation of the input layout to the camera’s 3D coordinate system has a key contribution to our method’s performance.
Table 2: Ablation study on our framework’s main components.
Ablation Setting FID $\downarrow$ KID $\downarrow$ SOC $\uparrow$
L-R 34.03 32.17 22.17
L-T 10.68 11.19 22.28
LACONIC 9.68 10.24 22.35
C.2 Baseline Comparison
We provide an additional qualitative comparison against the recent Build-A-Scene [7] method. While fundamentally different in its setting by adopting a training-free approach, the baseline also attempts to perform 3D layout guided image synthesis. To do so, objects are iteratively added to a given background by leveraging a depth-conditioned image generative model and by manipulating the attention maps of the pretrained backbone. Qualitative results from the same input bedroom layout are reported in Figure 10. Remarkably, our method more accurately reflects the input 3D structure, showcasing all and only the specified objects.
C.3 Perceptual Study
We conducted a perceptual study to compare our method with SceneCraft [77], evaluating both the overall quality of the generated images and their adherence to the input 3D layout and viewpoint.
Baseline Settings
Methods are compared on our custom bedroom dataset on which we train a SceneCraft model by optimizing jointly respective ControlNet [82] modules from depth and semantic maps. These conditioning inputs are rendered from the 3D layout and at the target camera view, as shown in Figure 11. We follow the official implementation and default parameters, using the same Stable Diffusion v1.5 backbone as our method. A qualitative comparison with this baseline is shown in Figure 14.
Study Design
We employed a two-alternative forced-choice (2AFC) test. For a given 3D layout and text prompt, participants were shown the pair of synthesized images and asked to choose the better one based on two separate criteria: (i) “Which image is the most realistic/natural?” and (ii) “Which image better respects the reference 3D layout and viewpoint?”. To prevent bias, the positions of the images were randomized for each trial. The study interface provided an interactive view of the 3D layout input and the ground truth image for reference, as shown in Figure 13.
Results
[…]
Figure 11: Conditioning inputs for the SceneCraft [77] baseline. The model is conditioned on depth (c) and segmentation (d) maps, which are rendered from the 3D bounding box layout (a) underlying a ground-truth bedroom image (b).
C.4 Qualitative Results
@@
We provide additional qualitative results on text-driven, layout-guided image synthesis in Figure LACONIC: A 3D Layout Adapter for Controllable Image Creation. These results highlight our method’s ability to adhere to the input semantic 3D layout while achieving out-of-distribution generalization to unseen concepts. Additionally, we report in Figure 15 and Figure 16 supplemental comparisons against the DM-FS and SceneCraft [77] baselines, both with and without providing an additional text condition. Our method consistently generates more realistic images with superior detail, 3D layout fidelity, and text prompt adherence. Motivating our 3D layout conditioning encoder, the DM-FS baseline appears to better respect the input 3D layout in comparison to SceneCraft [77]. We notice that both adapter-based approaches, ours and SceneCraft, leverage the pretrained T2I backbone, resulting in fewer artifacts in the generated images in comparison to DM-FS. Consequently, our method demonstrates a significant advantage in complex scenes with many objects, combining the expressiveness of our 3D layout encoder with the T2I backbone’s ability to generate detailed visual features.
Appendix D Limitations
Figure 12 illustrates failure cases and known limitations. Although our method demonstrates state-of-the-art 3D layout adherence, it can occasionally generate results with inconsistencies, such as missing objects, visual artifacts, or distorted perspectives.
@@ Appendix D Limitations
Figure 12: Failure Cases and Limitations. In (a), some objects, such as the windows, are missing in the generated image and the floor shape, while complex, is not accurately rendered. We can also observe visual stability issues, as shown in the result in (b), which features an inconsistent pattern on the wall resembling a misplaced frame above the bed. Finally, generated images can exhibit unnatural perspective, resulting in distorted floors and objects, as showcased in (c). This behavior may result from our assumption of consistent camera intrinsics across all data samples.
Appendix E Societal Impact
@@
Controllable generative AI for 3D environments has significant implications for both individuals and industries. We believe our method to predominantly yield positive societal impact. By enabling fine-grained control over generation, our approach makes 3D-aware content creation more intuitive and accessible, empowering not only artists and designers but also individuals with no prior expertise in 3D modeling. It also benefits industries that rely on realistic environmental rendering, such as urban planning, architecture, and digital twins for cities. A key advantage of our method is its parameter efficiency, which reduces the computational cost of training—an important factor given the high environmental impact of generative models [35, 28]. However, these benefits come with challenges. Since our approach does not natively incorporate recent advances in generative model watermarking [11, 60], it could be misused for potentially deceptive applications. Additionally, by significantly lowering the barrier to content generation, it may contribute to rapid job displacement in traditional 3D design fields. Another concern is the potential for bias propagation, as image generative models, including semantic and style-driven ones like ours, may unintentionally reinforce stereotypes [84]—for example, by associating certain colors with gender, by e.g., synthesizing a bedroom as blue for “boys” and pink for “girls”. Ensuring the responsible development and use of such technologies, while addressing ethical concerns and mitigating biases, is crucial.
Figure 13: Perceptual Study Interface. Users are prompted to independently select which generation result is (i) the most realistic and (ii) more in line with the input 3D layout and viewpoint.
Figure 14: Comparison with SceneCraft [77] on 3D layout-guided image synthesis. Our method demonstrates superior realism and adherence to the input 3D layout and viewpoint on our custom bedroom dataset.
@@
Figure 15: Additional 3D layout-guided image synthesis baseline comparisons (1/2). We can observe that our method produces more natural images that better respect the input 3D layout.